OpenStudy (anonymous):

write a parametrization for the equation of an ellipse at the origin crossing the x-axis at -3&3 and y-axis at -5&5

OpenStudy (anonymous):

{3 cos(t), 5 sin(t)}

OpenStudy (anonymous):

would you explain?

OpenStudy (anonymous):

since ellipse =x^2/a^2+y^2+b^2=1 since sin^2(t)+cos^2(t)=1 x^2/9+y^2/25=1 (x/3)^2+(y/5)^2=1 sin(t)=x/3 cos(t)=y/5 3sin(t)=x 5cos(t)=y 9sin^2(t)+25cos^2(t)=1 I believe thats a clear explaination

OpenStudy (anonymous):

i goofed on the x and y cos(t)=x/3 sin(t)=y/5 3cos(t)=x 5sin(t)=y 9cos^2(t)+25sin^2(t)=1